jim and sue have 72 coins that are nickels and dimes. The value of the coins is 5.80. How many of each do they have?
Answer by checkley77(12844)

N+D=72 OR N=72-D
.05(72-D)+.10D=5.80
3.6-.05D+.10D=5.80
.05D=5.80-3.60
.05D=2.20
D=2.20/.05
D=44 NUMBER OF DIMES.
N=72-44
N=28 NUMBER OF NICKELS.
PROOF:
.05*28+.10*44=5.80
1.40+4.40=5.80
5.80=5.80